OpenStudy (anonymous):

The Egan family paid $640 for auto insurance until ther 16-year-old son began driving the family car and the rates increased by 60%. Find the increased annual premium.

jimthompson5910 (jim_thompson5910):

640 + 60% of 640 = 640 + 0.6*640 = 640*1.6 = ??